OpenStudy (anonymous):

Simplify [7(x^2)(y^5) + 3(x^2)(y^5)] ÷ (5xy^2)

OpenStudy (anonymous):

We can see: \[ \frac{7(x^2)(y^5)+3(x^2)(y^5)}{5xy^2}=\frac{10x^2y^5}{5xy^2} \]And, using the exponent rule: \[ \frac{x^a}{x^b}=x^{a-b} \]We can solve for the rest.

OpenStudy (anonymous):

the answer will be 2xy^3..

OpenStudy (mayankdevnani):

\[7(x ^{2})(y ^{5})+3(x ^{2})(y ^{5})\div 5xy ^{2}\] \[7x ^{2}y ^{5}+3x ^{2}y ^{5}\div 5xy ^{2}\] \[7\div5 xy ^{3}+3\div5 xy ^{3}\] \[(7\div5+3\div5)xy ^{3}\] \[2xy ^{3}\]

OpenStudy (mayankdevnani):

\[\div\] means whole divided.....